OpenStudy (anonymous):

Simplify the following using exact values: (1+ (2/x)) / (1/x^2)

OpenStudy (anonymous):

As in \(\left(1+\frac{2}{x}\right)\bigg/\left(\frac{1}{x^2}\right)\) ?

OpenStudy (anonymous):

yes, exactly!